OpenStudy (anonymous):

What is the wavelength in Ångstroms of radiation used by an x-ray technician, with a frequency of 6.00 x 10^18 s-1? Please help me solve this problem by providing steps/work. The answer is .50X10^3 A. Thank you so much for helping.

OpenStudy (nincompoop):

were you given a formula in your lesson or class about wavelength, energy, etc?

OpenStudy (anonymous):

i believe it was C/F and C=3.0* 10^8

OpenStudy (astrophysics):

\[v = \lambda f\] is what you need, and yes x - rays go at the speed of light, so we are solving for \[\lambda = \frac{ v }{ f } = \frac{ c }{ f } = \frac{ 3 \times 10^8 m/s }{ 6 \times 10^{18} \frac{ 1 }{ s } } = 5 \times 10^{-11} m\] are you sure it's 10^18 Angstrom?

OpenStudy (astrophysics):

I meant Hz

OpenStudy (astrophysics):

To convert to angstrom just use \[1 ~ angstrom = 10^{-10}m\]

OpenStudy (astrophysics):

Are you there @gabz12

OpenStudy (anonymous):

i keep getting the 5*10^11 but the answer is 0.5 A

OpenStudy (astrophysics):

Are all the numbers in your question right? Your answer should be same as mine then but it's in meters, we have to convert it to angstrom as I just told you...

OpenStudy (astrophysics):

Yo nin do this

OpenStudy (astrophysics):

When I convert it to angstrom I get 0.5 A

OpenStudy (nincompoop):

I would solve it first using meters and then convert in the end

OpenStudy (astrophysics):

Aahah that's what I did

OpenStudy (astrophysics):

Oh I see he said the answer was 0.5 A

OpenStudy (astrophysics):

Ok so we have to convert it to angstrom \[\frac{ 5 \times 10^{-11}m }{ 10^{-10} m} \times 1 A = 0.5 A\] b00m

OpenStudy (astrophysics):

do you understand @gabz12

OpenStudy (nincompoop):

\(\huge \lambda = \frac{ v }{ f } = \frac{ c }{ f } = \frac{ 3.0 \times 10^8 ms^{-1} }{ 6.00 \times 10^{18} s^{-1} } \times \frac{1.0 A}{10^{-10}m}\) see how the units can be beautifully canceled out whole numbers simplified

OpenStudy (anonymous):

that formula should be provided right? ok i get it now thank you!

OpenStudy (astrophysics):

An easy way to think of this formula is, think about the units, notice how velocity is m/s, wavelength is m, and if you get frequency don't worry just think about the units and get 1/s. Another way notice how this formula IS exactly like \[v = \frac{ d }{ t }\] so that should give you a clue on what's going on.

OpenStudy (nincompoop):

OH ahaha wow I effed that one up

OpenStudy (nincompoop):

let me remove it